# The Illusion of Competence: Evaluating the Effect of Explanations on Users’ Mental Models of Visual Question Answering Systems

### v1

@@ Abstract
Judith Sieker, Simeon Junker, Ronja Utescher, Nazia Attari, Heiko Wersing, Hendrik Buschmeier, Sina Zarrieß Computational Linguistics Group, Linguistics Department, Bielefeld University Honda Research Institute Europe Digital Linguistics Lab, Linguistics Department, Bielefeld University
@@ 3 Approach
We conduct a study to investigate how users of an AI system perceive its limitations when it encounters tasks that it cannot perform perfectly. We aim to investigate whether providing explanations alongside model responses helps users build an appropriate mental model of the AI system’s capabilities and limitations. We control the AI system’s limitations by systematically manipulating its inputs. We design a questionnaire for users to judge specific aspects of the AI system’s capabilities. This allows us to measure whether users can diagnose which capabilities of the AI system have been perturbed through our explicit input manipulations. The design of our study is summarized in Figure 2 and will be explained in detail below.
VQA Task and Abilities
@@
Color vs. Grayscale Input
Our goal is to investigate whether explanations help users in diagnosing system limitations. To introduce these limitations in a controlled way, we manipulate the input of the VQA systems. Out of the four VQA capabilities explained above (color, shape, material, and scene), the color dimension lends itself to straightforward manipulation: during inference, systems either receive the image (i) in full color or (ii) in grayscale. This induced limitation resembles a situation where a multimodal AI model was trained on colored images but, at run-time, a camera/visual sensor is broken such that model inputs are perturbed. To make sure that this manipulation induces an incorrect model response, we only include items that are correctly answered with the full color image input but incorrectly answered with the grayscale image input. This item selection accounts for the fact that VQA models can be assumed to have further limitations that we cannot explicitly control for and exclude items (i) where the VQA does not generate the correct ground-truth answer for the colored image, and (ii) where the VQA generates the correct answer for the grayscale image. This gives us a clean set of items where the limitations of the AI system can be attributed to a particular error source. The participants in our study were unaware of the underlying color–grayscale manipulation: they saw images in color, along with the models’ answers and explanations. Our goal was to determine whether participants were able perceive the limitations of the model, i.e., whether they could identify the system’s lack of color recognition ability. See Figure 2 for an illustration of this set-up.
Experiments A and X
@@ 5 Results
All systems received high ratings in all competency and capability dimensions when tested in the color condition of Exp.A and X, on both datasets (see Table 9 in Appendix A.3). These ratings decreased in very similar ways in the grayscale condition. Therefore, we were able to use all items from all systems to test our hypotheses, generalizing over minor system differences. We discuss differences between datasets and models in Appendix A.3, since these were not essential for testing our hypotheses. Summaries of hypotheses and results are given in Table 2.
Hypotheses H1A and H1X
state our expectations on distinctions between the grayscale and color conditions in Exp.A and X, respectively. Figure 3 shows the distribution of participant ratings for the AI system’s ability to recognize colors, for the grayscale and color conditions in both experiments (see Figures 4, 5, 6, and 7 in Appendix A.3 for results on the other capabilities). In Exp.A and X, there is a consistent trend of better assessments when systems have been seen the color images compared to grayscale images, across different systems, datasets, and all capabilities. Most users rate the color capability with the highest rating in the color condition (Figure 3a/c) and with the lowest rating in the grayscale condition (Figure 3b/d). The same holds for all other capabilities and competency (Figures 4, 5, 6, and 7). This confirms hypothesis H1A, i.e., ratings for all capabilities decrease when the system does not see color. However, this does not support H1X, as we expected that only overall competence and capability to recognize colors would be rated lower in the grayscale condition when explanations were given, and not all capabilities. This suggests that the AI’s explanations did not help users diagnose the system’s limitation in the grayscale condition, as all capability dimensions are similarly affected in Exp.X.
@@
Hypotheses H2A and H2X
state our expectations for the grayscale condition. Table 1 presents the human evaluation results in Exp.A and X. Starting with Exp.A, Table 1 shows that all evaluation criteria in the grayscale condition receive relatively low scores. Interestingly, the manipulated capability, i.e., to recognize colors, does have slightly worse ratings than the other criteria (for most models and datasets). This outcome does not align with our expectation (H2A) as participants in Exp.A solely viewed the answers without access to explanations, making it difficult to discern which specific ability or (limitation) influenced the model’s answer. Results from Mann-Whitney U tests (see Table 4 in Appendix A.2) show significant differences between the ability to recognize colors and the ability to recognize other criteria for Exp.A (except for the models’ overall competence), contradicting hypothesis (H2A). This suggests that users in Exp.A were able to interpret incorrect system answers more than we expected.
@@
Hypothesis H3A/X
states our expectations regarding the differences between Exp.A and X for overall competency and color recognition ability.
@@ A.1 Materials Availability Statement
Our source code and the data from the human evaluation study will be made available in form of an accompanying data publication.
@@ A.3 Additional Results
Correlations between BERTscore and human judgements

### v2

@@ Abstract
Judith Sieker, Simeon Junker, Ronja Utescher, Nazia Attari, Heiko Wersing, Hendrik Buschmeier, Sina Zarrieß Computational Linguistics, Department of Linguistics, Bielefeld University Honda Research Institute Europe Digital Linguistics Lab, Department of Linguistics, Bielefeld University
@@ 3 Approach
We conduct a study to investigate how users of an AI system perceive its limitations when it encounters tasks that it cannot perform perfectly. We aim to investigate whether providing explanations alongside model responses helps users build an appropriate mental model of the AI system’s capabilities and limitations. In the following, we detail the rationale of our approach (Section 3.13.3) and the hypotheses of our experiments (Section 3.4).
3.1 How Can We Test Users’ Mental Models?
A key challenge for our study is that many modern AI systems are used in complex tasks that involve many interdependent capabilities simultaneously. This makes it difficult to isolate specific systems’ capabilities and to establish or control which limitations they have, even for the developers of a system. Indeed, common evaluation protocols in NLP mostly report overall system performance according to holistic metrics (e.g., accuracies) and rarely involve a detailed assessment of specific errors or capabilities (cf. van-miltenburg-etal-2021-underreporting). However, to assess whether users’ perception of a system is accurate, we need to have as much control as possible over its capabilities and limitations.
To address this challenge, our study adopts an experimental setting that simplifies some aspects of testing XAI in complex tasks. First, we focus on two well-studied VQA tasks, including a synthetic VQA task in which system capabilities are relatively easy to distinguish (Section 3.2). Second, to gain at least some control over the VQA systems’ limitations, we systematically manipulate their inputs in the dimension of color (Section 3.3). Third, we design a questionnaire for users to judge specific aspects of the system’s capabilities. This allows us to measure whether users can diagnose which capabilities of the system have been perturbed through our explicit input manipulations (Section 3.4). The design of our study is summarized in Figure 2 and will be explained in detail below.
3.2 VQA Task and Abilities
@@
3.3 Color vs. Grayscale Input
Our goal is to investigate whether explanations help users in diagnosing system limitations. To introduce these limitations in a controlled way, we manipulate the input of the VQA systems. Out of the four VQA capabilities explained above (color, shape, material, and scene), the color dimension lends itself to straightforward manipulation. During inference, systems either receive the image (i) in full color or (ii) in grayscale. This induced limitation resembles a situation where a multimodal AI model was trained on colored images but, at run-time, a camera/visual sensor is broken such that model inputs are perturbed. To ensure that this manipulation induces an incorrect model response, we only include items correctly answered with the full color image input but incorrectly answered with the grayscale image input. This item selection accounts for the fact that VQA models can be assumed to have further limitations that we cannot explicitly control for and exclude items (i) where the VQA does not generate the correct ground-truth answer for the colored image, and (ii) where the VQA generates the correct answer for the grayscale image. This gives us a clean set of items where the limitations of the AI system can be attributed to a particular error source. The participants in our study were unaware of the underlying color–grayscale manipulation: they saw images in color, along with the models’ answers and explanations. Our goal was to determine whether participants were able perceive the limitations of the model, i.e., whether they could identify the system’s lack of color recognition ability. See Figure 2 for an illustration of this set-up.
3.4 Experiments A and X
@@ 5 Results
All systems received high ratings in all competency and capability dimensions when tested in the color condition of Exp.A and X, on both datasets (see Table 9 in Appendix A.3). These ratings decreased in very similar ways in the grayscale condition. Therefore, we were able to use all items from all systems to test our hypotheses, generalizing over minor system differences. We discuss differences between datasets and models in Appendix A.3, since these were not essential for testing our hypotheses. See Table 2 for summaries of hypotheses and results.
5.1 Hypotheses H1A and H1X
Hypotheses H1A and H1X state our expectations on distinctions between the grayscale and color conditions in Exp.A and X, respectively. Figure 3 shows the distribution of participant ratings for the AI system’s ability to recognize colors, for the grayscale and color conditions in both experiments (see Figures 4, 5, 6, and 7 in Appendix A.3 for results on the other capabilities). In Exp.A and X, there is a consistent trend of better assessments when systems have been seen the color images compared to grayscale images, across different systems, datasets, and all capabilities. Most users rate the color capability with the highest rating in the color condition (Figure 3a/c) and with the lowest rating in the grayscale condition (Figure 3b/d). The same holds for all other capabilities and competency (Figures 4, 5, 6, and 7). This confirms hypothesis H1A, i.e., ratings for all capabilities decrease when the system does not see color. However, this does not support H1X, as we expected that only overall competence and capability to recognize colors would be rated lower in the grayscale condition when explanations were given, and not all capabilities. This suggests that the AI’s explanations did not help users diagnose the system’s limitation in the grayscale condition, as all capability dimensions are similarly affected in Exp.X.
@@ 5.1 Hypotheses H1A and H1X
5.2 Hypotheses H2A and H2X
Hypotheses H2A and H2X state our expectations for the grayscale condition. Table 1 presents the human evaluation results in Exp.A and X. Starting with Exp.A, Table 1 shows that all evaluation criteria in the grayscale condition receive relatively low scores. Interestingly, the manipulated capability, i.e., to recognize colors, does have slightly worse ratings than the other criteria (for most models and datasets). This outcome does not align with our expectation (H2A) as participants in Exp.A solely viewed the answers without access to explanations, making it difficult to discern which specific ability or (limitation) influenced the model’s answer. Results from Mann-Whitney U tests (see Table 4 in Appendix A.2) show significant differences between the ability to recognize colors and the ability to recognize other criteria for Exp.A (except for the models’ overall competence), contradicting hypothesis (H2A). This suggests that users in Exp.A were able to interpret incorrect system answers more than we expected.
@@ 5.2 Hypotheses H2A and H2X
5.3 Hypothesis H3A/X
Hypothesis H3A/X states our expectations regarding the differences between Exp.A and X for overall competency and color recognition ability.
@@ 7 Conclusion
Acknowledgments
The first author acknowledges financial support by the project “SAIL: SustAInable Life-cycle of Intelligent Socio-Technical Systems" (Grant ID NW21-059A), an initiative of the Ministry of Culture and Science of the State of North Rhine-Westphalia. We also acknowledge funding and support by the Deutsche Forschungsgemeinschaft (DFG) (TRR 318/1 2021 – 438445824) and Honda Research Institute Europe.
@@ A.1 Materials Availability Statement
The source code and data from our human evaluation study can be found at either of the following locations:
@@ A.3 Additional Results
Correlations between BERTscore and human judgments